OpenStudy (anonymous):

Simplify the following expression (2√4)^2

OpenStudy (anonymous):

@RoseDryer and @kausarsalley please help

OpenStudy (anonymous):

The \[\sqrt{4}=2\] So we have 2*2=4 4^2=16

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

No problem!